Project

General

Profile

11/03/14

The following notes were kindly taken by Dominick Rocco.

Attendance

Craig, Jon, Kanika, Ryan, Satish, Chris

Agenda

  • News (all)
  • FTS status (Dominick, all, Robert? )
  • Database troubles (Dominick, Jon?, Igor?)
  • Simulation generation status (Nate)
  • Calibration summary (Paola)
  • Reco FA14-10-28 status (Satish, Dominick )
  • Rapid turn around status (Dominick, Gavin, Evan)
  • Keep-up reco status (Dominick)
  • AOB

News

  • Craig wants to know how the FTS is doing

FTS

  • Things went great over the weekend for Dominick and Satish
  • Reached several hundred GB per hour on transfers

Database trouble

  • Overloaded the database web server with MC and data reco
  • Trying to reduce number of active connections, we seem to be saturating throughput
  • Chris wants an automatic email. Discussed this for a while, not everyone is convinced it would help
  • Dominick and Jon are working with Igor to test new configurations
  • Hopefully rapid progress will be made before Thursday's meeting

Simulation generation status (Nate)

  • Replacement ND files halfway done
  • Nate's priority is terrible, causing problems, projects died
  • Offsite this isn't a problem, but this is onsite
  • New geogittered files are going well, nonswap done, fluxwap and tauswap almost done

RHC files:

  • Need new flux files.
  • Waiting to hear from Raphael
  • Not reconstructing the ones we have since the flux is wrong.

Calibration summary (Paola)

  • Jobs are running, every day at 4am
  • FD cosmic data, runs 70677 and onward
  • 20 files consistently throwing errors ,
  • MC is all processed, as far as Paola knows

FA14-10-28 (Satish, Gavin, Dominick )

Satish, MC

Real-conditions MC:

  • Pretty much done
  • A handful of files failed, bad RunHistory information, email is out with details, waiting on experts
  • Jon says it's not missing info, something deeper in DBI

14 diblock ideal conditions MC

  • Too longer than expected
  • mostly done, small fraction failed, to be resubmitted

Dominick, data

  • Plagued by database trouble
  • Cosmics 90% finished, failures need to be re-run.
  • Numi is underway, pace is slow until database issues are sorted out

Gavin PID MC

  • Working on MC PD, going smoothly
  • Will start handing off lemsum datasets to Chris

Rapid turnaround (Dominick)

  • Gavin finished the nomask stuff
  • Dominick is waiting on new channel masks

Keep-up status (Dominick)

  • Still running manually because we're waiting to know if this round is sufficient
  • Chris will know in a week
  • If approved, will be handed off to SCD for automation

AOB

  • There was some talk about LEM prioritization
  • Gavin, Dominick and Chris will make sure Rapid-turnaround gets prioritized
  • Dataset definitions comments from Gavin
  • FCL definitions don't have nova.standard = True
  • Nate will fix them